OpenStudy (anonymous):

4x+5y=-18 x=1-4y use the substitution method to solve. Can someone please give me a walk through to solve this?

OpenStudy (mathteacher1729):

Didn't I answer this a few minutes ago?

OpenStudy (anonymous):

No lol....if you did it didnt show up on my page

OpenStudy (mathteacher1729):

No prob. In "Substitution" you get one variable all by itself and then you have x = (stuff) or y = (stuff). You then take that (stuff) and put it into the other equation wherever you see x or y.

OpenStudy (anonymous):

I got 4(1-4y)+5y=-18 4-16y+5y=-18 -5 -5 4-21y=-18 and from here I am lost?

OpenStudy (anonymous):

can you help me finish this?

OpenStudy (mathteacher1729):

You're right so far... 4 - 21y = -18 Now add 4 to both sides.

OpenStudy (anonymous):

OOOHHHHHH I got 16y+5=14 ??

OpenStudy (anonymous):

21y=-14 ?

OpenStudy (mathteacher1729):

Wait, no sorry. :( 4-16y+5y=-18 4 - 11y = -18 (negative 16 PLUS 5 = negative 11) Now subtract four to both sides. -11y = -22

OpenStudy (anonymous):

is y= -0.5?

OpenStudy (anonymous):

nevermind I think y=2 ?

OpenStudy (mathteacher1729):

Yup. y = 2. Now just go back to the 2nd equation and subsitute y = 2 to find out what x is. :)

OpenStudy (anonymous):

kk...does x=4?

OpenStudy (mathteacher1729):

How'd you get that?

OpenStudy (anonymous):

i had x=1-4y so I take x=1-4(2) and get x=2-8 divide and get 4 ?

OpenStudy (mathteacher1729):

Ah, ok. Unfortunately, that is not correct. FORTUNATELY, the answer is WAY EASIER than you've done. :D x = 1-4y. That's correct. x = 1-4(2) (x = one minus four times two) That's correct. x = 1 - 8 x = -7 Ta-daa! :D

OpenStudy (anonymous):

or negative 4

OpenStudy (anonymous):

OOH so I don't mu;ltiply the one by 2? I see. Thanks :)

OpenStudy (mathteacher1729):

Yup yup! You can check the graph of it by going to http://www.wolframalpha.com and typing plot 4x+5y=-18, x=1-4y